OpenStudy (anonymous):

Find the x-intercepts for the graph of the equation y=x^2-3x-4 The x-intercepts are??????????

OpenStudy (anonymous):

set y = 0 and solve

OpenStudy (anonymous):

\[x^2-3x-4=0\] \[(x-4)(x+1)=0\] \[x-4=0\] \[x=4\] \[x+1=0\] \[x=-1\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome